# The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning

### v3

@@ Appendix B Proofs for Results in Section 3
Consider first the case $\kappa=0$. Since $\kappa\geq\frac{f}{n-2f}$  [^5] Proposition 6, Allouah et al. (2023a), we have that $f=0$. Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$. Since $\kappa=0$, the output of $\mathbf{F}$ will correspond to the average of its input, which implies that
@@ Appendix C Proofs of the Results in Section 5
$\underbrace{\left\lVert{x_{1}}\right\rVert\leq\dots\leq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\leq\underbrace{\left\lVert{x_{k+2}}\right\rVert\leq\dots\leq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
@@ D.1 Datasets and Heterogeneity
In our experiments, we consider three standard image classification datasets, namely MNIST Deng (2012), Fashion-MNIST Xiao et al. (2017), and CIFAR-10 Krizhevsky et al. (2014). To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet Hsu et al. (2019a) distribution of parameter $\alpha$, as done in Hsu et al. (2019b); Allouah et al. (2023a); Farhadkhani et al. (2023). The smaller the $\alpha$, the more heterogeneous the setting. In our empirical evaluation, we set $\alpha\in\{0.1,0.5,1\}$ on (Fashion-)MNIST and $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$ on CIFAR-10 (refer to Figure 7). Furthermore, on MNIST and Fashion-MNIST, we also consider an extreme heterogeneity setting where the datapoints are sorted by increasing labels (0 to 9) and sequentially split equally among the honest workers.
@@ D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD Allouah et al. (2023a). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers. Furthermore, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD using a CNN of 1,310,922 parameters, in a distributed system comprising $n-f=16$ honest workers and $f=1$ adversarial worker. We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500. Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
@@
In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Allouah et al. (2023a).
L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25).
The comprehensive experimental setup, as well as the architecture of the models, are presented in Table 2.
@@
Honest workers $n-f=10$ $n-f=16$
Adversarial workers $f\in\{1,2,3,4,5,7,9\}$ $f=1$
Table 2: Experimental setup on (Fashion-)MNIST and CIFAR-10
@@ D.3 Byzantine Attacks
In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) Allen-Zhu et al. (2020), label-flipping (LF) Allen-Zhu et al. (2020), mimic Karimireddy et al. (2022), fall of empires (FOE) Xie et al. (2020), and a little is enough (ALIE) Baruch et al. (2019).
The exact functionality of the attacks is detailed below. In every step $t$, let $\overline{m}_{t}$ be an estimation of the true honest momentum at step $t$. In our experiments, we estimate $\overline{m}_{t}$ by averaging the momentums sent by the honest workers in step $t$ of Robust-DSHB. In other words, $\overline{m}_{t}=\frac{1}{|\mathcal{H}|}\sum\limits_{i\in\mathcal{H}}m_{t}^{(i)}$, where $m_{t}^{(i)}$ is the momentum computed by honest worker $w_{i}$ in step $t$.
@@
Since the FOE and ALIE attacks have $\tau$ as parameter, we implement in our experiments enhanced and adaptive versions of these attacks, where the attack factor is not constant and may be different in every iteration. In every step $t$, we determine the optimal attack factor $\tau_{t}$ through a grid search over a predefined range of values. More specifically, in every step $t$, $\tau_{t}$ takes the value that maximizes the damage inflicted by the adversarial workers, i.e., that maximizes the $l^{2}$ norm of the difference between the average of the honest momentums $\overline{m}_{t}$ and the output of the aggregation $R_{t}$ at the server.
@@ D.4 Benchmarking and Reproducibility
Appendix E Additional Experimental Results
In this section, we complete the experimental results that could not be placed in the main paper.
E.1 MNIST
(a) extreme
(b) $\alpha=0.5$
(c) $\alpha=0.1$
(d) $\alpha=0.1$
(e) $\alpha=1$
(f) $\alpha=1$
Figure 8: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers. We fix the heterogeneity level, and vary the the number of Byzantine workers $f$.
(a) $f=1$
(b) $f=3$
(c) $f=3$
(d) $f=5$
(e) $f=5$
Figure 9: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST with $10$ honest workers. We fix the number of Byzantine workers $f$, and vary the heterogeneity level.
(a) ALIE
(b) SF
(c) LF
(d) Mimic
Figure 10: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST (extreme heterogeneity) with $10$ honest workers and $f=1$, under several attacks. This complements Figure 3c of the main paper.
(a) FOE
(b) SF
(c) ALIE
(d) LF
(e) Mimic
Figure 11: Performance of Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed MNIST ($\alpha=0.1$) with $10$ honest workers and $f=3$, under several attacks.
(a) $f=3$
(b) $f=4$
(c) $f=5$
E.2 Fashion-MNIST
(a) $\alpha=0.1$
(b) $\alpha=0.1$
(c) $\alpha=0.05$
(d) $\alpha=0.05$
(e) extreme
(f) extreme
(a) $f=1$
(b) $f=1$
(c) $f=3$
(d) $f=3$
(e) $f=4$
(f) $f=4$
E.3 CIFAR-10
(a) $f=1$
(b) $f=1$
(a) ALIE with CWTM and GM
(b) FOE with CWMED and MK
(c) ALIE with CWMED and CWMED
Figure 16: Worst-case maximal accuracies achieved by Robust-DSGD when using ARC compared to no clipping, on heterogeneously-distributed CIFAR-10 ($\alpha=0.075$) with $16$ honest workers and $f=1$ executing ALIE and FOE. This figure complements Figure 5 in the main paper.
Appendix F Static Clipping vs ARC
@@
F.1 Limitations of Static Clipping
We perform our experiments using Robust-DSGD (see Algorithm 3) on (Fashion-)MNIST and CIFAR-10. We consider three different levels of heterogeneity, that we call moderate (corresponding to sampling from a Dirichlet distribution of parameter $\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are Byzantine. Furthermore, we train a convolutional neural network of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD on ResNet-18 He et al. (2015), in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine. We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500. Finally, we use the cross-entropy loss function with an $\ell_{2}$ regularization of $5\times 10^{-4}$.
In our experiments, we examine a wide range of static clipping parameters. Specifically, we choose $C\in\{0.02,0.2,2,20\}$ on MNIST.
(a) $f=3$
(b) $f=4$
Figure 17: Impact of the clipping strategy and heterogeneity level on the worst-case maximal accuracy achieved by Robust-DSGD against five Byzantine attacks on heterogeneous MNIST with $n=15$ workers. CWTM $\circ$ NNM is used as aggregation. DSGD reaches at least 98.5% in accuracy in all heterogeneity regimes.
@@ F.1 Limitations of Static Clipping
Unreliability due to the attack. First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process. To illustrate, while $C=2$ exhibits the best performance under the SF attack Allen-Zhu et al. (2020), depicted in Figure 1, it leads to a complete collapse of learning under LF Allen-Zhu et al. (2020). Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible). This highlights the fragility of static clipping, as the unpredictable nature of the Byzantine attack, a parameter that cannot be a priori known, can significantly degrade the performance of the chosen static clipping approach.
@@
Unreliability due to the heterogeneity model. Second, the robustness under static clipping is notably influenced by the level of heterogeneity present across the datasets of honest workers. To illustrate this impact, we present in Figure 17 the performances of Robust-DSGD for different levels of heterogeneity (moderate, high, and extreme) and different values of $C$. In the left plot of Figure 17, $C=2$ emerges as the best static clipping threshold when the heterogeneity is high whereas $C=20$ appears to be sub-optimal. On the other hand, under extreme heterogeneity, the accuracy associated to $C=2$ diminishes drastically while $C=20$ becomes a better choice. Intuitively, in heterogeneous scenarios, honest gradients become large in $l^{2}$-norm, due to the increasingly detrimental effect of Byzantine attacks. Consequently, we must increase the static clipping threshold. Failing to do so could introduce a significant bias. More details on this observation can be found in Figure 18. This highlights the intricate dependence of the performance of static clipping on data heterogeneity, and emphasizes the necessity to fine-tune static clipping strategies prior to the learning.
Figure 18: Distribution of the norms of honest gradients during the learning on heterogeneous MNIST. There are $f=3$ Byzantine workers (among $n=15$) executing the LF attack. CWTM $\circ$ NNM is used as aggregation method.
@@
Unreliability due to the number of Byzantine workers. Last but not least, the number of Byzantine workers $f$ also affects the efficacy of static clipping. As seen in Figure 17, $C=2$ leads to the highest accuracy among static clipping strategies when $f=3$ in high heterogeneity, but cannot be used when $f=4$ as its corresponding accuracy drops below 50% (see right plot of Figure 17).
@@
Overall, our empirical findings reveal that there is no single value of $C$ for which static clipping consistently delivers satisfactory performance across the various settings. We have also shown that the performance of static clipping is greatly influenced on typically uncontrollable parameters, such as data heterogeneity and the nature of the Byzantine attack. Indeed, it should be noted that explicitly estimating these parameters is challenging (if not impossible) in a distributed setting. First, since the server does not have direct access to the data, it cannot estimate the degree of heterogeneity. Second, as (by definition) a Byzantine worker can behave in an unpredictable manner, we cannot adjust $C$ to the attack(s) being executed by the Byzantine workers. This highlights the necessity for a robust clipping alternative that can naturally adapt to the setting in which it is deployed.
F.2 Empirical Benefits of ARC
@@
In contrast to static clipping, ARC is adaptive, delivering consistent performance across diverse levels of heterogeneity, Byzantine attacks, and number of Byzantine workers.
@@
The efficacy of our solution is illustrated in Figure 17, showcasing a consistently robust performance for all considered heterogeneity levels. Specifically, in scenarios of moderate heterogeneity where clipping may not be essential, ARC matches the performance of No clipping as well as static strategies $C=2$ and $C=20$. Conversely, the left plot of Figure 17 shows that under high and extreme heterogeneity, ARC surpasses the best static clipping strategy in terms of accuracy, highlighting the effectiveness of our approach.
@@
ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$. As depicted in Figures 1 and 17, ARC consistently yields robust performance, regardless of the Byzantine attack. In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 1). Furthermore, despite the increase in the number of Byzantine workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.
@@
Figure 19: Evolution of the adaptive clipping parameter $C_{t}$ of ARC compared to the norm of the honest mean during the learning on heterogeneous MNIST ($\alpha=1$). There are $f=3$ Byzantine workers among $n=15$. CWTM $\circ$ NNM is used as aggregation. ARC dynamically adjusts its clipping parameter $C_{t}$ based on the norms of honest momentums at step $t$, avoiding static over-clipping or under-clipping. This adaptability is evident in under all attacks, where $C_{t}$ consistently decreases with time, an expected behavior when reaching convergence.
(a) CWMed $\circ$ NNM
(b) GM $\circ$ NNM
(c) MK $\circ$ NNM

### v5

@@ Appendix B Proofs for Results in Section 3
Consider first the case $\kappa=0$. Since $\kappa\geq\frac{f}{n-2f}$  [^2] Proposition 6 (Allouah et al., 2023a), we have that $f=0$. Recall that we denote $y_{i}\coloneq\mathrm{clip}_{C}(x_{i})$ and $\overline{y}_{S}\coloneqq\frac{1}{\left\lvert{S}\right\rvert}\sum_{i\in S}y_{i}$. Since $\kappa=0$, the output of $\mathbf{F}$ will correspond to the average of its input, which implies that
@@ Appendix C Proofs of the Results in Section 5
$\underbrace{\left\lVert{x_{1}}\right\rVert\geq\dots\geq\left\lVert{x_{k+1}}\right\rVert}_{\text{at least $f+1$ vectors}}\geq\underbrace{\left\lVert{x_{k+2}}\right\rVert\geq\dots\geq\left\lVert{x_{n}}\right\rVert}_{\text{at most $n-f-1$ vectors}}$
@@ D.1 Datasets and Heterogeneity
In our experiments, we consider three standard image classification datasets, namely MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014). To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet distribution of parameter $\alpha$ (Hsu et al., 2019a), as done in (Hsu et al., 2019b; Allouah et al., 2023a; Farhadkhani et al., 2023). The smaller the $\alpha$, the more heterogeneous the setting. In our empirical evaluation, we set $\alpha\in\{0.1,0.5,1\}$ on (Fashion-)MNIST and $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$ on CIFAR-10 (refer to Figure 6). Furthermore, on MNIST and Fashion-MNIST, we also consider an extreme heterogeneity setting where the datapoints are sorted by increasing labels (0 to 9) and sequentially split equally among the honest workers.
@@ D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
We perform our experiments using Robust-DSGD, an order-optimal robust variant of Robust-DGD (Allouah et al., 2023a) (see Algorithm 3), in distributed systems of varying sizes. On MNIST and Fashion-MNIST, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we train a CNN of 1,310,922 parameters. We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500. Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$. In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Choffrut et al. (2024). L(#outputs) represents a fully-connected linear layer, R stands for ReLU activation, S stands for log-softmax, C(#channels) represents a fully-connected 2D-convolutional layer (kernel size 5, padding 0, stride 1), M stands for 2D-maxpool (kernel size 2), B stands for batch-normalization, and D represents dropout (with fixed probability 0.25). The comprehensive experimental setup and the architecture of the models are presented in Table 2.
@@
Table 2: Comprehensive experimental setup.
@@ D.3 Byzantine Attacks
In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) (Allen-Zhu et al., 2020), label-flipping (LF) (Allen-Zhu et al., 2020), mimic (Karimireddy et al., 2022), fall of empires (FOE) (Xie et al., 2020), and a little is enough (ALIE) (Baruch et al., 2019). The exact functionality of the attacks is detailed next. In every step $t$, let $\overline{m}_{t}$ be an estimation of the true honest momentum at step $t$. In our experiments, we estimate $\overline{m}_{t}$ by averaging the momentums sent by the honest workers in step $t$ of Robust-DSGD. In other words, $\overline{m}_{t}=\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}m_{t}^{(i)}$, where $m_{t}^{(i)}$ is the momentum computed by honest worker $w_{i}$ in step $t$.
@@
Since FOE and ALIE are parametrized by $\tau\in\mathbb{R}$, we implement in our experiments enhanced and adaptive versions of these attacks, where the attack factor is not constant and may be different in every iteration. In every step $t$, we determine the optimal attack factor $\tau_{t}$ through a grid search over a predefined range of values. More specifically, in every step $t$, $\tau_{t}$ takes the value that maximizes the damage inflicted by the adversarial workers, i.e., that maximizes the $\ell_{2}$-norm of the difference between the average of the honest momentums $\overline{m}_{t}$ and the output of the aggregation $R_{t}$ at the server.
@@ D.4 Benchmarking and Reproducibility
Appendix E ARC vs. Static Clipping
@@
E.1 Comparison with Static Clipping on MNIST
On MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are adversarial. We consider three different levels of heterogeneity: moderate ($\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D). In our experiments, we examine a wide range of static clipping parameters $C\in\{0.02,0.2,2,20\}$.
(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
Figure 7: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with CWTM $\circ$ NNM, under the SF and LF attacks. We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
@@ E.1 Comparison with Static Clipping on MNIST
Unreliability due to the attack. First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process. To illustrate, while $C=2$ exhibits the best performance under the SF attack (Allen-Zhu et al., 2020), depicted in Figure 7, it leads to a complete collapse of learning under LF (Allen-Zhu et al., 2020). Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible). This highlights the fragility of static clipping, as the unpredictable nature of the Byzantine attack, a parameter that cannot be a priori known, can significantly degrade the performance of the chosen static clipping approach.
@@
Unreliability due to the heterogeneity model. Second, the robustness under static clipping is notably influenced by the level of heterogeneity present across the datasets of honest workers. To illustrate this impact, we present in Figure 2 the performances of Robust-DSGD for different levels of heterogeneity (moderate, high, and extreme) and different values of $C$. In Figure 2(a), $C=2$ emerges as the best static clipping threshold when the heterogeneity is high whereas $C=20$ appears to be sub-optimal. On the other hand, under extreme heterogeneity, the accuracy associated to $C=2$ diminishes drastically while $C=20$ becomes a better choice. Intuitively, in heterogeneous scenarios, honest gradients become large in $\ell_{2}$-norm, due to the increasingly detrimental effect of Byzantine attacks. Consequently, we must increase the static clipping threshold. This highlights the intricate dependence of the performance of static clipping on data heterogeneity, and emphasizes the necessity to fine-tune static clipping strategies prior to the learning.
@@
Unreliability due to the number of adversarial workers. Last but not least, the number of adversarial workers $f$ also affects the efficacy of static clipping. As seen in Figure 2, $C=2$ leads to the highest accuracy among static clipping strategies when $f=3$ in high heterogeneity, but cannot be used when $f=4$ as its corresponding accuracy drops below 50% (see Figure 2(b)).
@@
Overall, our empirical findings reveal that there is no single value of $C$ for which static clipping consistently delivers satisfactory performance across the various settings. We have also shown that the performance of static clipping is greatly influenced on typically uncontrollable parameters, such as data heterogeneity and the nature of the Byzantine attack. Indeed, it should be noted that explicitly estimating these parameters is challenging (if not impossible) in a distributed setting. First, since the server does not have direct access to the data, it cannot estimate the degree of heterogeneity. Second, as (by definition) an adversarial worker can behave in an unpredictable manner, we cannot adjust $C$ to the attack(s) being executed by the adversarial workers. This highlights the necessity for a robust clipping alternative that can naturally adapt to the setting in which it is deployed.
@@
In contrast to static clipping, ARC is adaptive, delivering consistent performance across diverse levels of heterogeneity, Byzantine attacks, and number of adversarial workers.
@@
The efficacy of our solution is illustrated in Figure 2, showcasing a consistently robust performance for all considered heterogeneity levels. Specifically, in scenarios of moderate heterogeneity where clipping may not be essential, ARC matches the performance of No clipping as well as static strategies $C=2$ and $C=20$. Conversely, Figure 2(a) shows that under high and extreme heterogeneity, ARC surpasses the best static clipping strategy in terms of accuracy, highlighting the effectiveness of our approach.
@@
ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of adversarial workers $f$. As depicted in Figures 7 and 2, ARC consistently yields robust performance, regardless of the Byzantine attack. In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to Figure 7(b)). Furthermore, despite the increase in the number of adversarial workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.
@@
Figure 8: Evolution of the adaptive clipping parameter $C_{t}$ of ARC compared to the norm of the honest mean during the learning on heterogeneous MNIST ($\alpha=1$). There are $f=3$ adversarial workers among $n=15$. CWTM $\circ$ NNM is used as aggregation.